# Do We Really Need Permutations? Impact of Width Expansion on Linear Mode Connectivity

### v1

@@ D.1 Main Theorem
Definition D.3 (Sub-Gaussian random variable (Vershynin, 2018)).
@@
Definition D.4 (Sub-exponential random variable (Vershynin, 2018)).
@@
It is a classical fact on absolute moments of Gaussians (Haas, 2018) that
@@
By Theorem D.5, straightforward calculations yield
@@
Similarly, $\|{\sigma(u_{i}+v_{i})^{2}}\|_{\psi_{1}}\leq 32/3$ and $\|{(\sigma(u_{i})+\sigma(v_{i}))^{2}}\|_{\psi_{1}}\leq 32/3$. Hence each of these random variables is sub-exponential. Applying Bernstein’s inequality (Vershynin, 2018) with $K=32/3$, we obtain
@@
Appendix E Additional Experimental Results
E.1 CIFAR-100 Dataset
@@
E.2 Permutation-Based Method
@@
E.3 Random Permutations
Figure 13: Test accuracies of merged models with random permutations. When the model width is sufficiently large, the merged models perform well regardless of the random permutations applied.
For two models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ trained by SGD, we applied randomly chosen permutations $\pi_{a}$ and $\pi_{b}$ to each model, and merged the resulting models $\pi_{a}(\bm{\theta}_{a})$ and $\pi_{b}(\bm{\theta}_{b})$. The corresponding test accuracies are shown in Figure 13. It can be observed that, when the width is sufficiently large, high test accuracy is achieved regardless of the choice of permutations.
E.4 Standard Deviations of ReLU Inputs for All Layers
@@
Figure 14: Histogram of standard deviations around zero for each dimension of the ReLU inputs for all layers.
@@
In this section, we present histograms of the standard deviations of the ReLU inputs for all layers of each model. The results correspond to an MLP with width scaled by $16\times$, a VGG-11 scaled by $16\times$, and a ResNet-20 scaled by $32\times$. The histograms are shown in Figure 14. The vertical axis is plotted on a logarithmic scale. In addition, the values shown next to “Small input std.” indicate the proportion of input dimensions that fall into the leftmost bin. From Figure 14, it can be observed that in every layer, the vast majority of dimensions have very small standard deviations around zero.
E.5 Effect of Weight Decay on LEWC
(b) Average cosine similarity between $\sigma((\tilde{\bm{z}}_{\ell}^{(a)}+\tilde{\bm{z}}_{\ell}^{(b)})/2)$ and $(\sigma(\tilde{\bm{z}}_{\ell}^{(a)})+\sigma(\tilde{\bm{z}}_{\ell}^{(b)}))/2$, where $\tilde{\bm{z}}_{\ell}^{(a)}$ and $\tilde{\bm{z}}_{\ell}^{(b)}$ denote the pre-activations of the $\ell$-th layer of the two models.
Figure 15: Experimental results for models trained with a weaker weight decay of $10^{-4}$. From top to bottom, the figures present evaluations for LEWC, weak additivity of ReLU activations, and reciprocal orthogonality.
Figure 15 shows the experimental results when weight decay is reduced. Among the most prominent findings are those for VGG-11 and ResNet-20. First, Figure 15(a) indicates that LEWC is not satisfied in deeper layers when $\lambda=1/2$. Next, from Figure 15(b), we observe that the weak linearity of ReLU activations is also not satisfied in these models. Finally, Figure 15(c) demonstrates that reciprocal orthogonality does not hold either. Taken together, these results suggest that weakening the weight decay prevents LEWC and its two sufficient conditions from being satisfied.

### v2

@@ E.1 Main Theorem
Definition E.3 (Sub-Gaussian random variable (HDP_book)).
@@
Definition E.4 (Sub-exponential random variable (HDP_book)).
@@
It is a classical fact on absolute moments of Gaussians (Note_Absolute_Moments) that
@@
By Theorem E.5, straightforward calculations yield
@@
Similarly, $\|{\sigma(u_{i}+v_{i})^{2}}\|_{\psi_{1}}\leq 32/3$ and $\|{(\sigma(u_{i})+\sigma(v_{i}))^{2}}\|_{\psi_{1}}\leq 32/3$. Hence each of these random variables is sub-exponential. Applying Bernstein’s inequality (HDP_book) with $K=32/3$, we obtain
@@
Appendix F Additional Experimental Results
F.1 CIFAR-100 Dataset
@@
F.2 Permutation-Based Method
@@
F.3 Loss and Accuracy Barriers
(a) Accuracy barrier.
(b) Calibrated loss barrier.
Figure 17: Accuracy and calibrated loss barriers with and without the permutations identified by WM.
Figure 17 shows how the barrier changes with increasing width with and without the permutation found by WM. We observe that applying the permutation tends to reduce the barrier regardless of the width. However, when the width is sufficiently large, the barrier becomes nearly zero regardless of whether the permutation is used.
F.4 Random Permutations
(a) Test accuracies of merged models with random permutations.
(b) Difference in calibrated loss barriers between random permutations and the permutation identified by WM.
Figure 18: Effects of random permutations on accuracy and loss barriers.
For two models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ trained by SGD, we applied randomly chosen permutations $\pi_{a}$ and $\pi_{b}$ to each model, and merged the resulting models $\pi_{a}(\bm{\theta}_{a})$ and $\pi_{b}(\bm{\theta}_{b})$. The corresponding test accuracies are shown in Figure 18(a). We observe that, when the width is sufficiently large, high test accuracy is obtained regardless of the choice of permutations. Furthermore, Figure 18(b) shows the difference in calibrated loss barriers between using random permutations and using the permutation identified by WM. When the width is large enough, the barrier values remain nearly unchanged regardless of which permutation is applied.
F.5 Inverse Temperature
We present the inverse temperature values used for calibration when $\lambda=1/2$ in Figure 19. This inverse temperature corresponds to the factor by which the logits are divided during calibration. As the results indicate, under sufficiently wide conditions, the inverse temperature takes small values, compensating for the reduction in the logit norms caused by LEWC.
Figure 19: Inverse temperature used for logit calibration when $\lambda=1/2$.
F.6 Exponential Decay of Activation Norms
In Definition 4.1, we showed that the norm of the merged model’s output decreases exponentially. In this subsection, we experimentally verify whether this phenomenon actually occurs. The results are presented in Figure 20. For the MLP, as the depth increases, the activation norms of the merged model decrease exponentially, which is consistent with the prediction from LEWC. In contrast, for VGG-11 and ResNet-20, the norms do not decrease exponentially. This is presumably because batch normalization compensates for the variance.
Figure 20: The ratio of activation norms between the merged model and the original models, $\mathbb{E}\|f_{\ell}(\bm{x};(\bm{\theta}_{a}+\bm{\theta}_{b})/2)\|/\mathbb{E}\|f_{\ell}(\bm{x};\bm{\theta}_{a})+f_{\ell}(\bm{x};\bm{\theta}_{b})\|$, for $\lambda=1/2$. The test data are fed into the models.
F.7 Second Moments of ReLU Inputs for All Layers
@@
Figure 21: Histogram of the square root of the second (uncentered) moment for each dimension of the ReLU inputs across all layers.
@@
In this subsection, we present histograms of the square root of the second (uncentered) moment of the ReLU inputs for all layers of each model. The results correspond to an MLP with width scaled by $16\times$, a VGG-11 scaled by $16\times$, and a ResNet-20 scaled by $32\times$. The histograms are shown in Figure 21. The vertical axis is plotted on a logarithmic scale. In addition, the values shown next to “Small input std.” indicate the proportion of input dimensions that fall into the leftmost bin. From Figure 21, it can be observed that in every layer, the vast majority of dimensions have a very small second moment around zero.
F.8 Overlap Ratio of Active Neurons When Using Permutations
Figure 22: Proportion of input dimensions with a small square root of the second moment (“Small input moment”) and the proportion of non-overlapping large moment dimensions (“Non overlap”) between two trained models for the second ReLU layer, when using permutations found by WM.
Figure 22 shows the overlap ratio of active neurons in the second ReLU layer when applying the permutations identified by WM. This figure corresponds to Figure 6, where permutations are not used. When permutations are used, the non overlap ratio of active neurons clearly decreases, indicating that most neurons with large output second moments overlap between the two models. We attribute this to the fact that WM searches for permutations that reduce the $L^{2}$ distance between the weight matrices of the two models, which in turn tends to align their activation patterns.
F.9 Dependence on Weight Decay
(a) Accuracy barrier.
(b) Calibrated loss barrier.
Figure 23: Accuracy and loss barriers when varying the strength of weight decay.
Figure 23 shows the barrier values for permutation-free merging while varying the weight decay coefficient from 0 to 0.003. For sufficiently wide models, smaller weight decay results in smaller barriers.
F.10 Second Order Optimization
(a) Accuracy barrier.
(b) Calibrated loss barrier.
Figure 24: Accuracy and loss barriers with and without second-order optimization ($\lambda=1/2$).
(a) Accuracy.
(b) Calibrated loss.
Figure 25: Accuracy and loss of trained models with and without second order optimization.
In the main experiments, we used first order optimizers such as SGD and Adam. Here, we additionally examine the case where a second order optimization method is employed. Specifically, we adopt Sophia as a second order optimizer (Liu_ICLR_2024). The barrier results are shown in Figure 24. For reference, we also report the test accuracy and calibrated loss of the trained models in Figure 25. The training conditions, including hyperparameters, are kept the same as in the first order optimization setting. Therefore, the weight decay is fixed to 0.003.
From Figure 25, we observe that using a second order optimizer improves both accuracy and loss. However, the barriers are smaller when using first order optimizers. With second order optimization, the barriers remain clearly larger than zero for all models except for the MLP trained on MNIST. Nevertheless, the barrier decreases as the width increases, suggesting that LMC may hold when the width is sufficiently large. Since we did not retune hyperparameters for Sophia, it is also possible that using a larger weight decay or a different learning rate would make LMC more likely to emerge under second order optimization.
F.11 Effect of Skip Connections and Batch Normalization
(a) Accuracy barrier.
(b) Calibrated loss barrier.
Figure 26: Accuracy and calibrated loss barriers of models trained with and without skip connections or batch normalization.
(a) MLP with skip connections.
(b) MLP with batch normalization layers.
Figure 27: Ratio of mean norms $\mathbb{E}\|\bm{W}_{\ell}^{(a)}\bm{z}_{\ell}^{(b)}\|/\mathbb{E}\|\bm{W}_{\ell}^{(a)}\bm{z}_{\ell}^{(a)}\|$ for each layer, for MLPs with skip connections or batch normalization.
To clarify the effects of skip connections and batch normalization, we prepared two variants of the MLP: one with skip connections and one with batch normalization layers. For these models, we examined the barrier values and how easily reciprocal orthogonality is satisfied. Skip connections were added by summing each intermediate layer input with its output, for all hidden layers except the input and final layers. The resulting barrier values are shown in Figure 26, where the loss barrier is computed using calibrated loss. As shown in the figure, adding skip connections or batch normalization leads to smaller barriers.
To assess reciprocal orthogonality, we also evaluated the mean norm ratio $\mathbb{E}\|\bm{W}_{\ell}^{(a)}\bm{z}_{\ell}^{(b)}\|/\mathbb{E}\|\bm{W}_{\ell}^{(a)}\bm{z}_{\ell}^{(a)}\|$ for each layer, as reported in Figure 27. The results indicate that the presence or absence of skip connections or batch normalization does not substantially change this ratio.